OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). m + n = 1000 0.05m + 0.06n = 57

OpenStudy (anonymous):

m + n = 1000 time (-0.05) you have -0.05m-0.05n =-50 0.05m + 0.06n= 57 Add ----------------------- 0.01n = 7 n= 7/0.01 n=700

OpenStudy (anonymous):

substitution n=700 find m m + n = 1000 m + 700 = 1000 m= 1000-700 m = 300

OpenStudy (anonymous):

If u like my answer click good answer,ty

OpenStudy (anonymous):

what is x and y?

OpenStudy (anonymous):

x,y maybe m and n

OpenStudy (anonymous):

(x,y)=(300,700)

OpenStudy (anonymous):

answer in ordered pair form, (x, y). (x,y)=(300,700)